OpenStudy (anonymous):

Graph the function. Identify the vertex and axis of symmetry. f(x) = -3x^2 + 3x - 1

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \(-\frac{3}{2\times (-3)}=\frac{1}{2}\)

OpenStudy (anonymous):

this is also the "axis of symmetry" \(x=\frac{1}{2}\)